OpenStudy (anonymous):

Simplify. (aa^2)^3 / a^2 a^2

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

@nirmalnema

OpenStudy (anonymous):

\[\huge \frac{(aa^2)^3}{ a^2 a^2} = \frac{a^3 \times a^{2 \times 3}}{ a^2 \times a^2} \] \[\huge = \frac{a^3 \times a^{6}}{ a^2 \times a^2} = \frac {a^3 \times a^6}{a^4}\] \[\huge = \frac {a^3 \times a^6}{a^4}=a^3 \times a^2 = a^{3+2} = a^5\]

OpenStudy (anonymous):

@cancherolugano